### 13 Apr 2026 v2

# Retained-spin micropolar hydrodynamics from the Boltzmann–Curtiss equation: a generalized Chapman–Enskog construction

Abstract
We derive a retained-spin micropolar hydrodynamic closure from the Boltzmann–Curtiss equation using a generalized Chapman–Enskog construction in which the local mean spin is retained as a quasi-slow variable. Starting from the one-particle kinetic balance identities and the corresponding exact coarse-grained finite-size balances for mass, linear momentum, and intrinsic angular momentum, we keep the collisional-transfer contribution to the antisymmetric stress explicit in the spin balance, decompose the first-order source into irreducible scalar, axial, and symmetric-traceless sectors, and show explicitly how the standard micropolar constitutive structure with coefficients $(\eta,\xi,\eta_{r},\alpha,\beta,\gamma)$ emerges. This decomposition makes clear that the one-particle kinetic stress contributes only to the symmetric stress, whereas the rotational viscosity belongs to a collisional-transfer channel. For perfectly rough elastic hard spheres, we further obtain explicit dilute-gas estimates for the rotational viscosity $\eta_{r}$ from homogeneous spin relaxation and for the transverse spin-diffusion combination $\beta+\gamma$ from a transport-relaxation calculation. Targeted event-driven molecular-dynamics simulations are used as a posteriori checks: expanded homogeneous-spin density and roughness sweeps support the predicted $n^{2}$ and $K/(K+1)$ trends for $\eta_{r}$, while finite-$k$ transverse runs provide a qualitative diagnostic of the retained-spin response. The result is a self-contained derivation and coefficient-level estimate of retained-spin micropolar hydrodynamics that clarifies which parts of the closure are exact balance-law statements, which are first-order generalized Chapman–Enskog results, and which remain controlled rough-sphere estimates.
I Introduction
Micropolar and spin-carrying continuum theories extend classical fluid mechanics by promoting local micro-rotation to an independent field alongside the mass density, velocity, and temperature [7, 3, 9, 10, 11]. At the continuum level this introduces an antisymmetric stress channel, a couple stress, and a characteristic relaxation between vorticity and local spin. At the kinetic level the same physics is traced back to binary collisions that exchange translational and rotational angular momentum [5, 6, 18, 4, 16].
The basic links between the kinetic and continuum descriptions are classical, but in practice the derivation is scattered across several literatures. The exact moment balances are often stated without algebra. The Chapman–Enskog construction is typically described only in outline. The antisymmetric stress channel, which is the key to rotational viscosity, is especially susceptible to confusion because it is not generated by the symmetric one-particle kinetic stress and, for finite-size rough particles, must already be kept explicit in the exact coarse-grained intrinsic-spin balance. In addition, when a local spin field is retained at the hydrodynamic level, the calculation is no longer a strict hydrodynamic reduction in the narrow sense: the mean spin is not a collision invariant in general and must be treated as a quasi-slow variable in the sense of extended thermodynamics [19, 13].
The present derivation should be viewed against a literature that is real but highly dispersed. On the continuum side, early work on polar and spin-carrying fluids identified the antisymmetric stress and angular-momentum balance structure that later became standard in micropolar theory [7, 3, 9]. On the kinetic side, the Boltzmann–Curtiss and loaded-sphere programs introduced generalized kinetic equations with rotational degrees of freedom [5, 6], while rough-sphere transport and Chapman–Enskog analyses were developed for dilute gases [18, 4] and then extended to denser or more general polyatomic-fluid settings [16, 23, 24, 8]. These works provide many of the formal ingredients used below, but the relevant steps are scattered across continuum mechanics, molecular kinetic theory, and dense-fluid or rough-sphere transport theory rather than presented as a single, detailed retained-spin derivation.
Subsequent developments also proceeded largely along separate tracks. Micropolar continuum theory matured as an independent constitutive framework [10, 11]; granular and rough-sphere kinetic theories emphasized rapid granular flow, rough-particle transport, and hydrodynamic closures for particles with rotational degrees of freedom [14, 12, 17]; and recent Boltzmann–Curtiss-based studies have focused either on local-spin first-order closures or on higher-order constitutive models for polyatomic gases [26, 21, 15]. To our knowledge, however, a self-contained account that combines the exact balance-law structure, a retained-spin Chapman–Enskog construction, an irreducible first-order sector decomposition, and explicit dilute-gas rough-sphere estimates for the rotational coefficients has not been readily available in one place. The role of the present paper is therefore primarily reconstructive and expository: it collects these strands into a single notation and makes explicit which parts of the retained-spin closure are exact, which are formal Chapman–Enskog results, and which remain estimate-level rough-sphere evaluations.
The aim of the present paper is therefore narrow but useful. We do not attempt to develop a new response theory or a new application. The numerical material added here remains deliberately targeted rather than exhaustive: expanded EDMD benchmarks are used only to test the dilute rough-sphere estimates rather than to launch a separate transport-simulation program. We otherwise write out a detailed derivation of the retained-spin closure from the Boltzmann–Curtiss equation, with special attention to the exact balance-law structure, the irreducible tensor decomposition of the first-order source, the formal first-order coefficient problem, and the dilute-gas estimates of the key rotational coefficients for perfectly rough elastic hard spheres. The paper is intended to be readable line-by-line by a reader who wants to check the intermediate manipulations. A companion manuscript Ref. [25] addresses the response-theoretic consequences of the retained-spin closure, including EDMD observability and model discrimination among retained-spin, eliminated-spin, and polynomial higher-gradient descriptions. The present paper is instead derivation- and coefficient-estimate-oriented: its purpose is to provide the exact balance-law structure with the collisional-transfer torque channel made explicit, the generalized Chapman–Enskog construction with retained spin, and dilute-gas rough-sphere estimates for selected key rotational coefficients.
The scope is the following.
  1. (i)
    We derive the one-particle kinetic balance identities from the Boltzmann–Curtiss equation and write the corresponding exact finite-size intrinsic-spin balance with the collisional-transfer stress torque made explicit.
  2. (ii)
    We formulate a generalized Chapman–Enskog expansion retaining the local mean spin $\bm{\omega}_{0}$ as a quasi-slow variable.
  3. (iii)
    We decompose the first-order source into irreducible scalar, axial, and symmetric-traceless sectors and write the corresponding formal integral equations for the response functions.
  4. (iv)
    We show explicitly how the standard micropolar constitutive equations are recovered at first order, and we identify the precise place where the rotational viscosity $\eta_{r}$ enters.
  5. (v)
    For perfectly rough hard spheres, we derive explicit dilute-gas estimates of $\eta_{r}$ and of the transverse spin-diffusion combination $\beta+\gamma$.
  6. (vi)
    We supplement these rough-sphere estimates with targeted EDMD checks of homogeneous spin relaxation, including density and roughness sweeps, and of a finite-$k$ transverse retained-spin mode.
Two limitations should be stated at the outset. First, the present Chapman–Enskog construction is generalized or extended: the mean spin is treated as quasi-slow rather than as a strict collision invariant. More precisely, the retained-spin ordering adopted below assumes that the residual axial relaxation of the retained-spin manifold, represented in the collisional-transfer channel, is $O(\varepsilon)$, i.e. of the same asymptotic order as the first gradient corrections. Second, the antisymmetric stress channel is structurally closed at first order, but a fully coefficient-complete evaluation of $\eta_{r}$ for a rough-sphere collision operator would require an explicit collisional-transfer bracket. The rough-sphere formulas derived here should therefore be read as controlled low-density estimates rather than as the final word on the full antisymmetric Chapman–Enskog problem.
II Kinetic setting and macroscopic moments
II.1 State space and Boltzmann–Curtiss equation
We consider a dilute gas of rigid particles undergoing binary collisions. A single particle is characterized by its translational velocity $\bm{v}\in\mathbb{R}^{3}$ and intrinsic angular velocity $\bm{\omega}\in\mathbb{R}^{3}$. For perfectly rough spheres the orientation is irrelevant and the one-particle distribution is $f(\bm{x},\bm{v},\bm{\omega},t)$. For general rigid rotators an orientation $R\in\mathrm{SO}(3)$ must also be included and the distribution becomes $f(\bm{x},\bm{v},R,\bm{\omega},t)$. We write the phase-space measure as
$\mathrm{d}\Gamma=\mathrm{d}^{3}v\,\mathrm{d}^{3}\omega\quad\text{(rough spheres)},\qquad\mathrm{d}\Gamma=\mathrm{d}^{3}v\,\mathrm{d}R\,\mathrm{d}^{3}\omega\quad\text{(general rigid rotators)}.$ (1)
A convenient working form of the Boltzmann–Curtiss equation is
$\partial_{t}f+v_{j}\partial_{x_{j}}f+\frac{F^{\mathrm{ext}}_{j}}{m}\partial_{v_{j}}f+(\mathbb{I}^{-1}\bm{\tau}^{\mathrm{ext}})_{j}\partial_{\omega_{j}}f+\mathcal{R}[f]=C[f,f].$ (2)
Here $m$ is the particle mass, $\mathbb{I}$ is the inertia tensor, $\bm{F}^{\mathrm{ext}}$ is the external force on a particle, $\bm{\tau}^{\mathrm{ext}}$ is the external torque about the particle center, $\mathcal{R}[f]$ is the orientation-streaming term relevant for rigid rotators, and $C[f,f]$ is the binary collision operator. For perfectly rough spheres, $\mathcal{R}[f]=0$. We assume rapid decay of $f$ in $\bm{v}$ and $\bm{\omega}$ so that the integrations by parts below have no boundary contributions. The integration over $R\in\mathrm{SO}(3)$ contributes no boundary terms because the Haar measure is invariant and the manifold is compact.
II.2 Macroscopic fields and peculiar variables
The number density, mass density, mean velocity, and mean spin are defined by
$n(\bm{x},t)$ $:=\int f\,\mathrm{d}\Gamma,$ $\rho$ $:=mn,$ (3)
$u_{i}(\bm{x},t)$ $:=\frac{1}{n}\int v_{i}f\,\mathrm{d}\Gamma,$ $\omega_{0i}(\bm{x},t)$ $:=\frac{1}{n}\int\omega_{i}f\,\mathrm{d}\Gamma.$ (4)
The associated peculiar variables are
$c_{i}:=v_{i}-u_{i},\qquad\Omega_{i}:=\omega_{i}-\omega_{0i}.$ (5)
By construction,
$\int c_{i}f\,\mathrm{d}\Gamma=0,\qquad\int\Omega_{i}f\,\mathrm{d}\Gamma=0.$ (6)
From this point onward we restrict attention to isotropic microinertia,
$\mathbb{I}=I\,\mathrm{Id},\qquad J:=\frac{I}{m},$ (7)
which covers perfectly rough spheres and isotropically distributed rigid rotators.
For the local quasi-equilibrium introduced later we use a single temperature $T$. The full heat-flux problem is not developed in this paper; the focus is on the mechanical and spin transport channels.
II.3 Collision invariants
For elastic binary collisions the standard collision invariants are the particle number, linear momentum, total angular momentum, and total kinetic energy [2, 1]. In particular we assume that
$\int C[f,f]\,\mathrm{d}\Gamma$ $=0,$ (8)
$\int mv_{i}\,C[f,f]\,\mathrm{d}\Gamma$ $=0,$ (9)
$\int\Bigl[\varepsilon_{ijk}x_{j}mv_{k}+(\mathbb{I}\bm{\omega})_{i}\Bigr]C[f,f]\,\mathrm{d}\Gamma$ $=0.$ (10)
The total-energy invariant will not be used explicitly below and is therefore omitted from the displayed list. Equations (8)–(10) are sufficient for the one-particle moment identities derived below. For finite-size rough particles, however, the exact local coarse-grained momentum and angular-momentum balances contain additional contact-transfer fluxes. In particular, the antisymmetric part of the total stress produces a local intrinsic-spin torque. We therefore distinguish below between the one-particle kinetic moments obtained directly from Eq. (2) and the exact finite-size balance laws used for the retained-spin micropolar closure.
III Exact balance laws from the kinetic equation
III.1 Mass balance
Integrating Eq. (2) over phase space gives
$\int\partial_{t}f\,\mathrm{d}\Gamma+\int v_{j}\partial_{x_{j}}f\,\mathrm{d}\Gamma+\int\frac{F^{\mathrm{ext}}_{j}}{m}\partial_{v_{j}}f\,\mathrm{d}\Gamma+\int(\mathbb{I}^{-1}\bm{\tau}^{\mathrm{ext}})_{j}\partial_{\omega_{j}}f\,\mathrm{d}\Gamma+\int\mathcal{R}[f]\,\mathrm{d}\Gamma=\int C[f,f]\,\mathrm{d}\Gamma.$ (11)
Each term is evaluated separately.
$\int\partial_{t}f\,\mathrm{d}\Gamma$ $=\partial_{t}\int f\,\mathrm{d}\Gamma=\partial_{t}n,$ (12)
$\int v_{j}\partial_{x_{j}}f\,\mathrm{d}\Gamma$ $=\partial_{x_{j}}\int v_{j}f\,\mathrm{d}\Gamma=\partial_{x_{j}}(nu_{j}),$ (13)
$\int\frac{F^{\mathrm{ext}}_{j}}{m}\partial_{v_{j}}f\,\mathrm{d}\Gamma$ $=\frac{F^{\mathrm{ext}}_{j}}{m}\int\partial_{v_{j}}f\,\mathrm{d}\Gamma=0,$ (14)
$\int(\mathbb{I}^{-1}\bm{\tau}^{\mathrm{ext}})_{j}\partial_{\omega_{j}}f\,\mathrm{d}\Gamma$ $=(\mathbb{I}^{-1}\bm{\tau}^{\mathrm{ext}})_{j}\int\partial_{\omega_{j}}f\,\mathrm{d}\Gamma=0,$ (15)
$\int\mathcal{R}[f]\,\mathrm{d}\Gamma$ $=0,$ (16)
$\int C[f,f]\,\mathrm{d}\Gamma$ $=0.$ (17)
Thus
$\partial_{t}n+\partial_{x_{j}}(nu_{j})=0.$ (18)
Multiplying by $m$ gives the continuity equation
$\partial_{t}\rho+\bm{\nabla}\!\cdot(\rho\bm{u})=0.$ (19)
III.2 Linear-momentum balance
Multiply Eq. (2) by $mv_{i}$ and integrate over phase space. The time and transport terms are
$\int mv_{i}\partial_{t}f\,\mathrm{d}\Gamma$ $=\partial_{t}\int mv_{i}f\,\mathrm{d}\Gamma=\partial_{t}(\rho u_{i}),$ (20)
$\int mv_{i}v_{j}\partial_{x_{j}}f\,\mathrm{d}\Gamma$ $=\partial_{x_{j}}\int mv_{i}v_{j}f\,\mathrm{d}\Gamma.$ (21)
For the force term we integrate by parts in velocity:
$\int mv_{i}\frac{F^{\mathrm{ext}}_{j}}{m}\partial_{v_{j}}f\,\mathrm{d}\Gamma$ $=F^{\mathrm{ext}}_{j}\int v_{i}\partial_{v_{j}}f\,\mathrm{d}\Gamma$
$=F^{\mathrm{ext}}_{j}\int\Bigl[\partial_{v_{j}}(v_{i}f)-\delta_{ij}f\Bigr]\,\mathrm{d}\Gamma$
$=-F^{\mathrm{ext}}_{i}\int f\,\mathrm{d}\Gamma=-nF^{\mathrm{ext}}_{i}.$ (22)
The torque term vanishes because $v_{i}$ is independent of $\bm{\omega}$:
$\int mv_{i}(\mathbb{I}^{-1}\bm{\tau}^{\mathrm{ext}})_{j}\partial_{\omega_{j}}f\,\mathrm{d}\Gamma=0.$ (23)
The orientation-streaming term also vanishes,
$\int mv_{i}\mathcal{R}[f]\,\mathrm{d}\Gamma=0,$ (24)
and the collision term vanishes by Eq. (9):
$\int mv_{i}C[f,f]\,\mathrm{d}\Gamma=0.$ (25)
We therefore obtain
$\partial_{t}(\rho u_{i})+\partial_{x_{j}}\Pi_{ij}^{(k)}=\rho F_{i},\qquad F_{i}:=\frac{F_{i}^{\mathrm{ext}}}{m},$ (26)
where the kinetic momentum-flux tensor is
$\Pi_{ij}^{(k)}:=\int mv_{i}v_{j}f\,\mathrm{d}\Gamma.$ (27)
To separate convection from stress, write $v_{i}=u_{i}+c_{i}$ and expand:
$\Pi_{ij}^{(k)}$ $=\int m(u_{i}+c_{i})(u_{j}+c_{j})f\,\mathrm{d}\Gamma$
$=mu_{i}u_{j}\int f\,\mathrm{d}\Gamma+mu_{i}\int c_{j}f\,\mathrm{d}\Gamma+mu_{j}\int c_{i}f\,\mathrm{d}\Gamma+\int mc_{i}c_{j}f\,\mathrm{d}\Gamma.$ (28)
Using Eq. (6) gives
$\Pi_{ij}^{(k)}=\rho u_{i}u_{j}+\int mc_{i}c_{j}f\,\mathrm{d}\Gamma.$ (29)
The kinetic Cauchy stress is therefore
$\sigma_{ij}^{(k)}:=-\int mc_{i}c_{j}f\,\mathrm{d}\Gamma.$ (30)
Because $c_{i}c_{j}$ is symmetric, $\sigma^{(k)}_{ij}$ is symmetric. Equation (26) may then be written as
$\partial_{t}(\rho u_{i})+\partial_{x_{j}}(\rho u_{i}u_{j}-\sigma_{ij}^{(k)})=\rho F_{i},$ (31)
or, equivalently,
$\rho\frac{\mathrm{D}u_{i}}{\mathrm{D}t}=\partial_{x_{j}}\sigma^{(k)}_{ij}+\rho F_{i}.$ (32)
For structured or finite-size particles, Eq. (32) identifies only the one-particle kinetic contribution. The exact coarse-grained local momentum balance contains an additional contact-transfer stress $\sigma_{ij}^{(\mathrm{int})}$. We therefore reserve the notation
$\sigma_{ij}:=\sigma_{ij}^{(k)}+\sigma_{ij}^{(\mathrm{int})}$ (33)
for the total Cauchy stress and define the corresponding total momentum flux by
$\Pi_{ij}:=\rho u_{i}u_{j}-\sigma_{ij}=\Pi_{ij}^{(k)}-\sigma_{ij}^{(\mathrm{int})}.$ (34)
The exact finite-size momentum balance is then
$\partial_{t}(\rho u_{i})+\partial_{x_{j}}\Pi_{ij}=\rho F_{i},\qquad\rho\frac{\mathrm{D}u_{i}}{\mathrm{D}t}=\partial_{x_{j}}\sigma_{ij}+\rho F_{i}.$ (35)
In a point-particle theory $\sigma_{ij}^{(\mathrm{int})}\equiv 0$ and Eq. (35) reduces to Eq. (32).
III.3 Total and intrinsic angular momentum
The exact intrinsic-spin balance is obtained most cleanly by first writing the balance of total angular momentum and then subtracting the orbital part. For a single particle, the total angular momentum about the origin is
$\ell_{i}:=\varepsilon_{ijk}x_{j}mv_{k}+(\mathbb{I}\bm{\omega})_{i}.$ (36)
Multiply Eq. (2) by $\ell_{i}$ and integrate over phase space.
The time-derivative term gives
$\int\ell_{i}\partial_{t}f\,\mathrm{d}\Gamma=\partial_{t}\int\ell_{i}f\,\mathrm{d}\Gamma.$ (37)
For the transport term we must differentiate $\ell_{i}$ with respect to $\bm{x}$ because $x_{j}$ is not integrated over:
$\int\ell_{i}v_{j}\partial_{x_{j}}f\,\mathrm{d}\Gamma$ $=\partial_{x_{j}}\int\ell_{i}v_{j}f\,\mathrm{d}\Gamma-\int(\partial_{x_{j}}\ell_{i})v_{j}f\,\mathrm{d}\Gamma.$ (38)
Now
$\partial_{x_{j}}\ell_{i}=\partial_{x_{j}}(\varepsilon_{ipq}x_{p}mv_{q})=\varepsilon_{ipq}\delta_{jp}mv_{q}=\varepsilon_{ijq}mv_{q},$ (39)
so Eq. (38) becomes
$\int\ell_{i}v_{j}\partial_{x_{j}}f\,\mathrm{d}\Gamma=\partial_{x_{j}}\int\ell_{i}v_{j}f\,\mathrm{d}\Gamma-\varepsilon_{ijq}\int mv_{q}v_{j}f\,\mathrm{d}\Gamma.$ (40)
The force term gives the expected external moment:
$\int\ell_{i}\frac{F_{j}^{\mathrm{ext}}}{m}\partial_{v_{j}}f\,\mathrm{d}\Gamma$ $=-\int\partial_{v_{j}}\ell_{i}\,\frac{F_{j}^{\mathrm{ext}}}{m}f\,\mathrm{d}\Gamma$
$=-\varepsilon_{ipq}x_{p}F_{q}^{\mathrm{ext}}\int f\,\mathrm{d}\Gamma$
$=-\rho(\bm{x}\times\bm{F})_{i}.$ (41)
The torque term is
$\int\ell_{i}(\mathbb{I}^{-1}\bm{\tau}^{\mathrm{ext}})_{j}\partial_{\omega_{j}}f\,\mathrm{d}\Gamma$ $=-\int\partial_{\omega_{j}}\ell_{i}\,(\mathbb{I}^{-1}\bm{\tau}^{\mathrm{ext}})_{j}f\,\mathrm{d}\Gamma$
$=-\tau_{i}^{\mathrm{ext}}\int f\,\mathrm{d}\Gamma=-n\tau_{i}^{\mathrm{ext}}.$ (42)
The orientation-streaming term integrates to zero, and the collision term vanishes because $\ell_{i}$ is a collision invariant:
$\int\ell_{i}C[f,f]\,\mathrm{d}\Gamma=0.$ (43)
Collecting terms yields the balance of total angular momentum,
$\partial_{t}\mathcal{L}_{i}+\partial_{x_{j}}\mathcal{J}_{ij}=\varepsilon_{ijq}\Pi^{(k)}_{qj}+\rho(\bm{x}\times\bm{F})_{i}+n\tau_{i}^{\mathrm{ext}},$ (44)
with
$\mathcal{L}_{i}:=\int\ell_{i}f\,\mathrm{d}\Gamma,\qquad\mathcal{J}_{ij}:=\int\ell_{i}v_{j}f\,\mathrm{d}\Gamma.$ (45)
Now split total angular momentum into orbital and intrinsic parts. Since
$\int mv_{q}f\,\mathrm{d}\Gamma=\rho u_{q},\qquad\int(\mathbb{I}\bm{\omega})_{i}f\,\mathrm{d}\Gamma=In\omega_{0i}=\rho J\omega_{0i},$ (46)
we have
$\mathcal{L}_{i}=\varepsilon_{ipq}x_{p}\rho u_{q}+\rho J\omega_{0i}.$ (47)
Likewise, the one-particle part of the angular-momentum flux is
$\mathcal{J}^{(k)}_{ij}$ $=\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj}+\int(\mathbb{I}\bm{\omega})_{i}v_{j}f\,\mathrm{d}\Gamma$
$=\varepsilon_{ipq}x_{p}\Pi^{(k)}_{qj}+\rho J\omega_{0i}u_{j}-m_{ij}^{(k)},$ (48)
where the kinetic couple stress is defined as minus the nonconvective intrinsic flux
$m_{ij}^{(k)}:=-\int(\mathbb{I}\bm{\omega})_{i}(v_{j}-u_{j})f\,\mathrm{d}\Gamma=-\int(\mathbb{I}\bm{\omega})_{i}c_{j}f\,\mathrm{d}\Gamma.$ (49)
For finite-size particles the exact coarse-grained total angular-momentum flux contains additional contact-transfer contributions. We therefore write
$\mathcal{J}_{ij}:=\varepsilon_{ipq}x_{p}\Pi_{qj}+\rho J\omega_{0i}u_{j}-m_{ij},$ (50)
where $m_{ij}$ is the exact intrinsic flux. In the dilute first-order closure developed below, its leading contribution is represented by the one-particle moment $m_{ij}^{(k)}$.
Next derive the balance of orbital angular momentum directly from the exact finite-size momentum equation (35). Multiplying by $\varepsilon_{ipq}x_{p}$ gives
$\partial_{t}(\varepsilon_{ipq}x_{p}\rho u_{q})+\varepsilon_{ipq}x_{p}\partial_{x_{j}}\Pi_{qj}$ $=\rho(\bm{x}\times\bm{F})_{i}.$ (51)
Using the product rule,
$\varepsilon_{ipq}x_{p}\partial_{x_{j}}\Pi_{qj}=\partial_{x_{j}}(\varepsilon_{ipq}x_{p}\Pi_{qj})-\varepsilon_{ijq}\Pi_{qj},$ (52)
so the orbital balance is
$\partial_{t}(\varepsilon_{ipq}x_{p}\rho u_{q})+\partial_{x_{j}}(\varepsilon_{ipq}x_{p}\Pi_{qj})=\rho(\bm{x}\times\bm{F})_{i}+\varepsilon_{ijq}\Pi_{qj}.$ (53)
The corresponding exact coarse-grained balance of total angular momentum reads
$\partial_{t}\mathcal{L}_{i}+\partial_{x_{j}}\mathcal{J}_{ij}=\rho(\bm{x}\times\bm{F})_{i}+\rho G_{i},$ (54)
where, as before,
$\rho G_{i}:=n\tau_{i}^{\mathrm{ext}}.$ (55)
Subtracting Eq. (53) from Eq. (54) and using Eqs. (47) and (50) gives the exact finite-size intrinsic-spin balance
$\partial_{t}(\rho J\omega_{0i})+\partial_{x_{j}}(\rho J\omega_{0i}u_{j}-m_{ij})=-\varepsilon_{ijq}\Pi_{qj}+\rho G_{i}.$ (56)
Since the convective part $\rho u_{q}u_{j}$ is symmetric, $\varepsilon_{ijq}\Pi_{qj}=-\varepsilon_{ijk}\sigma_{jk}$, and we may write
$\partial_{t}(\rho J\bm{\omega}_{0})+\bm{\nabla}\!\cdot(\rho J\bm{\omega}_{0}\otimes\bm{u}-\bm{m})=\bm{\tau}^{(\sigma)}+\rho\bm{G},\qquad\tau_{i}^{(\sigma)}:=-\varepsilon_{ijk}\sigma_{jk}.$ (57)
Using the continuity equation in the usual way gives the nonconservative form
$\rho J\frac{\mathrm{D}\omega_{0i}}{\mathrm{D}t}=\partial_{x_{j}}m_{ij}-\varepsilon_{ijk}\sigma_{jk}+\rho G_{i}.$ (58)
III.4 Where does the antisymmetric stress appear?
Equation (58) already contains the antisymmetric-stress torque through the exact total stress. In a point-particle theory $\sigma_{ij}=\sigma_{ij}^{(k)}$ is symmetric, so $\tau_{i}^{(\sigma)}=0$ and the intrinsic balance reduces to the purely kinetic flux form. For finite-size rough particles, however, the contact-transfer part $\sigma_{ij}^{(\mathrm{int})}$ need not be symmetric, and
$\tau_{i}^{(\sigma)}:=-\varepsilon_{ijk}\sigma_{jk}=-\varepsilon_{ijk}\sigma_{jk}^{(\mathrm{int})}$ (59)
represents the local orbital–intrinsic exchange. This is the exact channel that later carries the rotational viscosity $\eta_{r}$; it is not an additional constitutive term to be appended after the fact. We return to this point in Sec. V.
IV Generalized Chapman–Enskog construction with retained spin
IV.1 Scaling and extended-hydrodynamic viewpoint
Let $L$ be a macroscopic length scale and $\ell$ a mean free path. The Knudsen number is
$\varepsilon:=\frac{\ell}{L}\ll 1.$ (60)
The Chapman–Enskog method assumes that the distribution function depends on $(\bm{x},t)$ only through slowly varying macroscopic fields and expands
$f=f^{(0)}+\varepsilon f^{(1)}+\varepsilon^{2}f^{(2)}+\cdots,\qquad\partial_{t}=\partial_{t}^{(0)}+\varepsilon\partial_{t}^{(1)}+\cdots.$ (61)
The retained-spin closure uses the fields
$(n,\bm{u},T,\bm{\omega}_{0})$ (62)
as the coordinates of the reference manifold.
Strictly speaking, only the collision invariants are guaranteed to remain slow in the ordinary dilute-gas hydrodynamic limit. The mean spin is generally not a strict invariant and relaxes on a collisional time scale. Retaining $\bm{\omega}_{0}$ therefore corresponds to a generalized or extended-hydrodynamic description [19, 13]. This viewpoint is not a mathematical defect but the intended regime: we want a closure that keeps the local spin explicit rather than eliminating it instantaneously.
IV.2 Local quasi-equilibrium distribution
The natural reference state for the retained-spin theory is the maximum-entropy distribution consistent with $(n,\bm{u},T,\bm{\omega}_{0})$, namely
$f^{(0)}=n\left(\frac{m}{2\pi k_{B}T}\right)^{3/2}\left(\frac{I}{2\pi k_{B}T}\right)^{3/2}\exp\!\left[-\frac{mc^{2}}{2k_{B}T}-\frac{I\Omega^{2}}{2k_{B}T}\right].$ (63)
For general isotropic rotators the second prefactor becomes $(\det\mathbb{I})^{1/2}/(2\pi k_{B}T)^{3/2}$, but the scalar form (63) is sufficient for the present paper.
The zeroth-order moments are immediate. The kinetic stress is isotropic,
$\sigma_{ij}^{(k,0)}=-\int mc_{i}c_{j}f^{(0)}\,\mathrm{d}\Gamma=-P\,\delta_{ij},\qquad P=nk_{B}T,$ (64)
and the couple stress vanishes,
$m_{ij}^{(k,0)}=-\int I\omega_{i}c_{j}f^{(0)}\,\mathrm{d}\Gamma=0,$ (65)
because $c_{j}$ has zero mean under the local Maxwellian. Viscous and spin-diffusive effects therefore arise only at first order.
At this point the extended-hydrodynamic ordering should be stated explicitly. For a strict classical Chapman–Enskog construction, the local Maxwellian lies on the exact null manifold of the one-particle collision operator, so that $C[f^{(0)},f^{(0)}]=0$. In the retained-spin closure for finite-size rough particles, however, the axial collisional-transfer channel associated with the exact spin torque need not vanish identically on the retained-spin manifold. We therefore introduce an auxiliary bookkeeping parameter $\delta_{s}$ that measures the residual axial relaxation strength and write the matched-order axial source symbolically as
$\delta_{s}=O(\varepsilon),\qquad\mathcal{S}^{(\mathrm{ax})}_{i}a_{i}=\mathcal{S}^{(\mathrm{ax},\nabla)}_{i}a_{i}+\delta_{s}\,\mathcal{R}^{(\mathrm{ax})}[f^{(0)}].$
Here $\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]$ denotes the residual axial contribution of the retained-spin manifold, understood to belong to the collisional-transfer channel rather than to the symmetric one-particle kinetic stress. It is orthogonal to the strict mass, linear-momentum, and energy invariants,
$\int\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0,\qquad\int mv_{i}\,\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0,\qquad\int\left(\frac{m}{2}v^{2}+\frac{I}{2}\omega^{2}\right)\mathcal{R}^{(\mathrm{ax})}[f^{(0)}]\,\mathrm{d}\Gamma=0.$
Operationally, the present retained-spin closure may therefore be regarded as a double expansion in $\varepsilon$ and $\delta_{s}$, with the bookkeeping $\delta_{s}\sim\varepsilon$. Under this matched ordering, the residual axial relaxation enters at the same asymptotic order as the usual gradient-driven first-order source. This is the precise sense in which $\bm{\omega}_{0}$ is treated as a quasi-slow variable in the present paper. If instead $\delta_{s}=O(1)$, the irreducible sector decomposition derived below still identifies the correct constitutive channels, but the construction should then be read as a formal extended-hydrodynamic closure rather than as a strict one-parameter Chapman–Enskog limit. The explicit rough-sphere estimates derived later are therefore coefficient evaluations within the same retained-spin constitutive structure; they are not, by themselves, an additional proof of the ordering $\delta_{s}\sim\varepsilon$.
IV.3 Linearized first-order equation
Write the first-order correction in the standard multiplicative form
$f^{(1)}=f^{(0)}\phi.$ (66)
The Chapman–Enskog matching conditions keep the retained fields entirely in $f^{(0)}$; in particular,
$\int f^{(1)}\,\mathrm{d}\Gamma=0,\qquad\int c_{i}f^{(1)}\,\mathrm{d}\Gamma=0,\qquad\int\Omega_{i}f^{(1)}\,\mathrm{d}\Gamma=0.$ (67)
Define the linearized collision operator around $f^{(0)}$ by
$\mathcal{L}[\phi]:=-\frac{1}{f^{(0)}}\Bigl(C[f^{(0)}\phi,f^{(0)}]+C[f^{(0)},f^{(0)}\phi]\Bigr).$ (68)
Then the first-order Chapman–Enskog equation is
$\mathcal{L}[\phi]=\mathcal{S},$ (69)
where the source $\mathcal{S}$ collects the $O(\varepsilon)$ streaming contributions together with the matched-order residual axial collisional-transfer contribution described above. The standard density, temperature, and heat-flux sectors follow the classical Chapman–Enskog pattern [2, 22]. Here we focus on the sectors relevant to the stress and couple stress.
IV.4 Useful derivatives of the local Maxwellian
The logarithm of (63) is
$\ln f^{(0)}=\ln n-3\ln(2\pi k_{B}T)+\frac{3}{2}\ln m+\frac{3}{2}\ln I-\frac{mc^{2}}{2k_{B}T}-\frac{I\Omega^{2}}{2k_{B}T}.$ (70)
At fixed $(\bm{v},\bm{\omega})$,
$\partial_{x_{\ell}}c_{i}=-\partial_{x_{\ell}}u_{i},\qquad\partial_{x_{\ell}}\Omega_{i}=-\partial_{x_{\ell}}\omega_{0i}.$ (71)
Consequently,
$\partial_{x_{\ell}}c^{2}$ $=2c_{i}\partial_{x_{\ell}}c_{i}=-2c_{i}\partial_{x_{\ell}}u_{i},$ (72)
$\partial_{x_{\ell}}\Omega^{2}$ $=2\Omega_{i}\partial_{x_{\ell}}\Omega_{i}=-2\Omega_{i}\partial_{x_{\ell}}\omega_{0i}.$ (73)
Differentiating $\ln f^{(0)}$ with respect to $x_{\ell}$ gives
$\partial_{x_{\ell}}\ln f^{(0)}$ $=\partial_{x_{\ell}}\ln n-3\partial_{x_{\ell}}\ln T+\frac{mc^{2}+I\Omega^{2}}{2k_{B}T}\partial_{x_{\ell}}\ln T+\frac{m}{k_{B}T}c_{i}\partial_{x_{\ell}}u_{i}+\frac{I}{k_{B}T}\Omega_{i}\partial_{x_{\ell}}\omega_{0i}$
$=\partial_{x_{\ell}}\ln n+\left[\frac{mc^{2}+I\Omega^{2}}{2k_{B}T}-3\right]\partial_{x_{\ell}}\ln T+\frac{m}{k_{B}T}c_{i}\partial_{x_{\ell}}u_{i}+\frac{I}{k_{B}T}\Omega_{i}\partial_{x_{\ell}}\omega_{0i}.$ (74)
This is the basic starting point for the first-order source decomposition.
The full source $\mathcal{S}$ also contains Euler-level time derivatives of the slow fields. After the usual elimination of the invariant channels by the zeroth-order balance equations, the mechanical and spin-gradient part of the source takes the schematic form
$\mathcal{S}_{\mathrm{mech}}=-\frac{m}{k_{B}T}c_{i}c_{j}\partial_{j}u_{i}-\frac{I}{k_{B}T}c_{j}\Omega_{i}\partial_{j}\omega_{0i}+\mathcal{S}^{(\mathrm{ax})}_{i}a_{i}+\cdots,$ (75)
where the dots denote density, temperature, and heat-flux channels, and
$a_{i}:=\frac{1}{2}\zeta_{i}-\omega_{0i},\qquad\zeta_{i}:=(\bm{\nabla}\!\times\bm{u})_{i}.$ (76)
The last term $\mathcal{S}^{(\mathrm{ax})}_{i}a_{i}$ is written separately because the axial mismatch sector is not generated by the symmetric one-particle kinetic stress alone. In the matched ordering described above, this symbol also absorbs the $O(\varepsilon)$ residual axial relaxation of the retained-spin quasi-equilibrium manifold. The axial sector therefore belongs to the collisional-transfer channel and will be treated formally in Sec. VI.
IV.5 Irreducible decomposition of the velocity and spin gradients
Introduce the standard decomposition of the velocity gradient,
$\partial_{i}u_{j}=D_{ij}+\frac{1}{3}\theta\,\delta_{ij}+W_{ij},\qquad\theta:=\partial_{k}u_{k},\qquad W_{ij}:=\frac{1}{2}(\partial_{i}u_{j}-\partial_{j}u_{i}),$ (77)
where $D_{ij}$ is symmetric traceless:
$D_{ij}=\frac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\frac{1}{3}\theta\,\delta_{ij}.$ (78)
Because $c_{i}c_{j}$ is symmetric,
$c_{i}c_{j}W_{ij}=0.$ (79)
Therefore
$c_{i}c_{j}\partial_{j}u_{i}=\left\langle c_{i}c_{j}\right\rangle D_{ij}+\frac{1}{3}c^{2}\theta.$ (80)
Equation (80) is crucial: the one-particle kinetic streaming source contains no axial contribution proportional to the antisymmetric velocity gradient.
Now decompose the spin gradient into irreducible parts. Define
$q:=\partial_{k}\omega_{0k},\qquad b_{i}:=\frac{1}{2}\varepsilon_{ijk}\partial_{j}\omega_{0k}=\frac{1}{2}(\bm{\nabla}\!\times\bm{\omega}_{0})_{i},$ (81)
and the symmetric traceless tensor
$E_{ij}:=\frac{1}{2}(\partial_{i}\omega_{0j}+\partial_{j}\omega_{0i})-\frac{1}{3}q\,\delta_{ij}.$ (82)
Then
$\partial_{i}\omega_{0j}=E_{ij}+\frac{1}{3}q\,\delta_{ij}+\varepsilon_{ijk}b_{k}.$ (83)
Using this in the second term of Eq. (75) gives
$c_{j}\Omega_{i}\partial_{j}\omega_{0i}$ $=c_{j}\Omega_{i}\left(E_{ji}+\frac{1}{3}q\,\delta_{ji}+\varepsilon_{jik}b_{k}\right)$
$=\left\langle c_{j}\Omega_{i}\right\rangle E_{ij}+\frac{1}{3}(\bm{c}\cdot\bm{\Omega})q+(\bm{c}\times\bm{\Omega})_{k}b_{k}.$ (84)
The natural basis functions in the spin-gradient sector are therefore
$Y^{(0)}:=\bm{c}\cdot\bm{\Omega},\qquad Y_{i}^{(1)}:=(\bm{c}\times\bm{\Omega})_{i},\qquad Y_{ij}^{(2)}:=\left\langle c_{j}\Omega_{i}\right\rangle.$ (85)
V First-order constitutive structure
V.1 Sectorwise expansion of the first-order correction
Rotational invariance implies that the first-order correction can be expanded in irreducible sectors. Restricting attention to the channels relevant for the stress and couple stress, we write
$\phi=B^{(\mathrm{dev})}_{ij}D_{ij}+B^{(\mathrm{tr})}\theta+X_{i}a_{i}+Y\,q+Z_{i}b_{i}+Z_{ij}E_{ij}+\cdots,$ (86)
where $Z_{ij}$ is symmetric traceless in $(i,j)$ and the omitted terms belong to heat and diffusion sectors. The response functions $B^{(\mathrm{dev})}_{ij},B^{(\mathrm{tr})},X_{i},Y,Z_{i},Z_{ij}$ are functions of the microscopic variables $(\bm{c},\bm{\Omega})$.
Because the linearized collision operator commutes with proper rotations in the isotropic case, the distinct irreducible sectors decouple at first order. This is the basic reason why one may speak of the shear channel, the bulk channel, the spin-diffusion channels, and the axial mismatch channel.
V.2 Total stress at first order
The first-order kinetic stress is
$\sigma_{ij}^{(k,1)}=-\int mc_{i}c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (87)
The kernel $c_{i}c_{j}f^{(0)}$ is symmetric in $(i,j)$ and even under $\bm{\Omega}\mapsto-\bm{\Omega}$. The axial-mismatch and spin-gradient sectors in Eq. (86) are odd in $\bm{\Omega}$ and therefore integrate to zero against this kernel; in particular, no $q\,\delta_{ij}$ or $E_{ij}$ contribution survives in the kinetic stress. Thus only the symmetric velocity-gradient sectors contribute to $\sigma_{ij}^{(k,1)}$, and rotational invariance forces the result to be of the form
$\sigma_{ij}^{(k)}=-P\,\delta_{ij}+2\eta D_{ij}+\xi\theta\,\delta_{ij}.$ (88)
The coefficients $\eta$ and $\xi$ are the shear and bulk viscosities.
The antisymmetric channel belongs to the collisional-transfer part of the total stress. Introduce the axial mismatch tensor
$A_{ij}:=W_{ij}-\varepsilon_{ijk}\omega_{0k}.$ (89)
Since
$W_{ij}=\frac{1}{2}\varepsilon_{ijk}\zeta_{k},$ (90)
we can also write
$A_{ij}=\varepsilon_{ijk}a_{k},\qquad a_{k}=\frac{1}{2}\zeta_{k}-\omega_{0k}.$ (91)
Isotropy and parity imply that the first-order intrinsic stress in this channel must itself be antisymmetric and proportional to $A_{ij}$, so we define
$\sigma_{ij}^{(\mathrm{int},1)}=-2\eta_{r}A_{ij}.$ (92)
Equivalently,
$\sigma_{ij}^{(\mathrm{int},1)}=\eta_{r}\varepsilon_{ijk}(2\omega_{0k}-\zeta_{k}),\qquad-\varepsilon_{ijk}\sigma_{jk}^{(\mathrm{int},1)}=2\eta_{r}(\zeta_{i}-2\omega_{0i}).$ (93)
Combining Eqs. (88) and (92), the total first-order stress is
$\sigma_{ij}=-P\,\delta_{ij}+2\eta D_{ij}+\xi\theta\,\delta_{ij}-2\eta_{r}W_{ij}+2\eta_{r}\varepsilon_{ijk}\omega_{0k}.$ (94)
This is the standard micropolar stress with the present sign convention.
V.3 Divergence of the total stress
We now compute $\partial_{j}\sigma_{ij}$ explicitly because the intermediate vector identities are often skipped. Using Eq. (94),
$\partial_{j}\sigma_{ij}$ $=-\partial_{i}P+2\eta\,\partial_{j}D_{ij}+\xi\,\partial_{i}\theta-2\eta_{r}\,\partial_{j}W_{ij}+2\eta_{r}\,\partial_{j}(\varepsilon_{ijk}\omega_{0k}).$ (95)
The divergence of the symmetric traceless rate is
$\partial_{j}D_{ij}$ $=\partial_{j}\left[\frac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\frac{1}{3}\theta\,\delta_{ij}\right]$
$=\frac{1}{2}\partial_{i}\partial_{j}u_{j}+\frac{1}{2}\partial_{j}\partial_{j}u_{i}-\frac{1}{3}\partial_{i}\theta$
$=\frac{1}{2}\partial_{i}\theta+\frac{1}{2}\bm{\nabla}^{2}u_{i}-\frac{1}{3}\partial_{i}\theta$
$=\frac{1}{2}\bm{\nabla}^{2}u_{i}+\frac{1}{6}\partial_{i}\theta.$ (96)
Hence
$2\eta\,\partial_{j}D_{ij}=\eta\bm{\nabla}^{2}u_{i}+\frac{\eta}{3}\partial_{i}\theta.$ (97)
For the antisymmetric velocity-gradient term,
$-2\eta_{r}\partial_{j}W_{ij}$ $=-\eta_{r}\partial_{j}(\partial_{i}u_{j}-\partial_{j}u_{i})$
$=-\eta_{r}(\partial_{i}\theta-\bm{\nabla}^{2}u_{i})$
$=\eta_{r}\bm{\nabla}^{2}u_{i}-\eta_{r}\partial_{i}\theta.$ (98)
Finally,
$2\eta_{r}\partial_{j}(\varepsilon_{ijk}\omega_{0k})=2\eta_{r}(\bm{\nabla}\!\times\bm{\omega}_{0})_{i}.$ (99)
Substituting Eqs. (97), (98), and (99) into Eq. (95) gives
$\partial_{j}\sigma_{ij}=-\partial_{i}P+(\eta+\eta_{r})\bm{\nabla}^{2}u_{i}+\left(\frac{\eta}{3}+\xi-\eta_{r}\right)\partial_{i}\theta+2\eta_{r}(\bm{\nabla}\!\times\bm{\omega}_{0})_{i}.$ (100)
The momentum equation therefore becomes
$\rho\frac{\mathrm{D}u_{i}}{\mathrm{D}t}=-\partial_{i}P+(\eta+\eta_{r})\bm{\nabla}^{2}u_{i}+\left(\frac{\eta}{3}+\xi-\eta_{r}\right)\partial_{i}\theta+2\eta_{r}(\bm{\nabla}\!\times\bm{\omega}_{0})_{i}+\rho F_{i}.$ (101)
V.4 Couple stress at first order
The leading one-particle contribution to the first-order couple stress is obtained from Eq. (49):
$m_{ij}^{(k,1)}=-\int I\omega_{i}c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (102)
Writing $\omega_{i}=\Omega_{i}+\omega_{0i}$ gives
$m_{ij}^{(k,1)}=-\int I\Omega_{i}c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma-\omega_{0i}\int Ic_{j}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (103)
The second term vanishes by the Chapman–Enskog matching condition $\int c_{j}f^{(1)}\,\mathrm{d}\Gamma=\int c_{j}f^{(0)}\phi\,\mathrm{d}\Gamma=0$ from Eq. (67). Introduce the microscopic kernel
$Q_{ij}:=I\Omega_{i}c_{j}.$ (104)
Then
$m_{ij}^{(k,1)}=-\int Q_{ij}f^{(0)}\phi\,\mathrm{d}\Gamma.$ (105)
At dilute first order we represent the total couple stress by this leading one-particle contribution, so only the $q$, $b_{i}$, and $E_{ij}$ sectors contribute. Rotational invariance therefore implies the general form
$m_{ij}=\lambda_{0}q\,\delta_{ij}+\lambda_{1}\varepsilon_{ijk}b_{k}+\lambda_{2}E_{ij}.$ (106)
The coefficients are linear functionals of the three response functions $Y,Z_{i},Z_{ij}$. A convenient extraction is
$\lambda_{0}$ $=-\frac{1}{3}\int Q_{kk}f^{(0)}Y\,\mathrm{d}\Gamma,$ (107)
$\lambda_{1}$ $=-\frac{1}{6}\varepsilon_{ijk}\int Q_{ij}f^{(0)}Z_{k}\,\mathrm{d}\Gamma,$ (108)
$\lambda_{2}$ $=-\frac{1}{5}\int\left\langle Q_{ij}\right\rangle f^{(0)}Z_{ij}\,\mathrm{d}\Gamma.$ (109)
To compare with the conventional micropolar notation, start from
$m_{ij}=\alpha\,\delta_{ij}\,\partial_{k}\omega_{0k}+\beta(\partial_{i}\omega_{0j}+\partial_{j}\omega_{0i})+\gamma(\partial_{j}\omega_{0i}-\partial_{i}\omega_{0j}).$ (110)
Using Eq. (83),
$\partial_{i}\omega_{0j}+\partial_{j}\omega_{0i}$ $=2E_{ij}+\frac{2}{3}q\,\delta_{ij},$ (111)
$\partial_{j}\omega_{0i}-\partial_{i}\omega_{0j}$ $=-2\varepsilon_{ijk}b_{k}.$ (112)
Substituting into Eq. (110) gives
$m_{ij}=\left(\alpha+\frac{2}{3}\beta\right)q\,\delta_{ij}-2\gamma\,\varepsilon_{ijk}b_{k}+2\beta E_{ij}.$ (113)
Comparing with Eq. (106) yields
$\lambda_{0}=\alpha+\frac{2}{3}\beta,\qquad\lambda_{1}=-2\gamma,\qquad\lambda_{2}=2\beta.$ (114)
Equivalently,
$\beta=\frac{1}{2}\lambda_{2},\qquad\gamma=-\frac{1}{2}\lambda_{1},\qquad\alpha=\lambda_{0}-\frac{1}{3}\lambda_{2}.$ (115)
V.5 Divergence of the couple stress
Starting from Eq. (110),
$\partial_{j}m_{ij}$ $=\alpha\,\partial_{i}\partial_{k}\omega_{0k}+\beta\,\partial_{j}(\partial_{i}\omega_{0j}+\partial_{j}\omega_{0i})+\gamma\,\partial_{j}(\partial_{j}\omega_{0i}-\partial_{i}\omega_{0j})$
$=\alpha\,\partial_{i}q+\beta(\partial_{i}q+\bm{\nabla}^{2}\omega_{0i})+\gamma(\bm{\nabla}^{2}\omega_{0i}-\partial_{i}q)$
$=(\beta+\gamma)\bm{\nabla}^{2}\omega_{0i}+(\alpha+\beta-\gamma)\partial_{i}q.$ (116)
The exact spin balance (58) therefore becomes, after inserting the first-order constitutive forms for $m_{ij}$ and $\sigma_{ij}$,
$\rho J\frac{\mathrm{D}\omega_{0i}}{\mathrm{D}t}=(\beta+\gamma)\bm{\nabla}^{2}\omega_{0i}+(\alpha+\beta-\gamma)\partial_{i}(\bm{\nabla}\!\cdot\bm{\omega}_{0})+2\eta_{r}(\zeta_{i}-2\omega_{0i})+\rho G_{i}.$ (117)
Equations (101) and (117) are the retained-spin micropolar equations derived from the first-order constitutive structure.
VI Formal first-order coefficient problem
The derivation above determines the tensorial form of the constitutive laws. We now spell out the first-order Chapman–Enskog problems that define the coefficients themselves.
VI.1 Shear and bulk sectors
The shear source in Eq. (75) is proportional to $\left\langle c_{i}c_{j}\right\rangle D_{ij}$, so the shear response satisfies a tensor equation of the form
$\mathcal{L}[B^{(\mathrm{dev})}_{ij}]=-\frac{m}{k_{B}T}\left\langle c_{i}c_{j}\right\rangle.$ (118)
The shear viscosity is then extracted from
$\eta=-\frac{1}{10}\int m\left\langle c_{i}c_{j}\right\rangle f^{(0)}B^{(\mathrm{dev})}_{ij}\,\mathrm{d}\Gamma.$ (119)
Likewise the scalar bulk channel may be written as
$\mathcal{L}[B^{(\mathrm{tr})}]=S^{(\mathrm{tr})},$ (120)
with the corresponding moment formula
$\xi=-\frac{1}{3}\int m\left(\frac{1}{3}c^{2}-\frac{k_{B}T}{m}\right)f^{(0)}B^{(\mathrm{tr})}\,\mathrm{d}\Gamma.$ (121)
The precise scalar basis in the bulk channel depends on how the invariant part is projected out. This is standard and is not the focus of the present paper.
VI.2 Spin-diffusion sectors
From Eqs. (75) and (84), the three spin-gradient channels satisfy
$\mathcal{L}[Y]$ $=-\frac{I}{3k_{B}T}(\bm{c}\cdot\bm{\Omega}),$ (122)
$\mathcal{L}[Z_{i}]$ $=-\frac{I}{k_{B}T}(\bm{c}\times\bm{\Omega})_{i},$ (123)
$\mathcal{L}[Z_{ij}]$ $=-\frac{I}{k_{B}T}\left\langle c_{j}\Omega_{i}\right\rangle.$ (124)
Together with Eqs. (107)–(109), these equations define $\lambda_{0}$, $\lambda_{1}$, and $\lambda_{2}$, hence $(\alpha,\beta,\gamma)$ through Eq. (115).
At this point the formal closure of the spin-diffusion sector is complete in the structural sense: the irreducible driving fields, the linear equations, and the moment formulas are all explicit.
VI.3 Rotational-viscosity sector
The rotational-viscosity sector is subtler. As Eq. (79) showed, the symmetric kinetic stress kernel $mc_{i}c_{j}$ cannot generate an axial stress. The coefficient $\eta_{r}$ therefore belongs to the collisional-transfer part of the total stress. Denote the corresponding linear stress functional by $\mathfrak{S}_{ij}[\phi]$. In the present axial channel this gives
$\sigma_{ij}^{(\mathrm{int},1)}=\mathfrak{S}_{ij}[X_{k}a_{k}].$ (125)
By isotropy, parity, and linearity this map must take the form
$\mathfrak{S}_{ij}[X_{k}a_{k}]=-2\eta_{r}A_{ij}.$ (126)
The corresponding response function satisfies a linear equation
$\mathcal{L}[X_{i}]=\mathcal{S}^{(\mathrm{ax})}_{i}$, (127)
where $\mathcal{S}^{(\mathrm{ax})}_{i}$ denotes the net axial source produced by the collisional-transfer mechanism and by the retained-spin extended manifold.
Equations (126) and (127) are the formal first-order definition of $\eta_{r}$. To evaluate $\eta_{r}$ for a specific rough-sphere collision model one still needs an explicit collisional-transfer representation for $\mathfrak{S}_{ij}$ and the associated collision bracket. This is precisely the point at which the antisymmetric problem becomes harder than the ordinary symmetric stress problem.
VI.4 What is closed and what is not
It is worth isolating the logical status of the first-order calculation.
  • The one-particle balance identities are fully derived, and the exact finite-size antisymmetric-stress torque channel is made explicit.
  • The first-order constitutive form is fully derived.
  • The spin-diffusion coefficient problem is formally closed by Eqs. (122)–(124) and (107)–(109).
  • The rotational-viscosity channel is structurally closed, but its coefficient-level evaluation still requires an explicit collisional-transfer bracket.
This is the reason for presenting the rough-sphere formulas in the next section as controlled dilute-gas estimates.
VII Perfectly rough hard spheres
VII.1 Collision rule
We now specialize to identical perfectly rough elastic hard spheres of diameter $a$, mass $m$, and moment of inertia $I$. Introduce the standard reduced moment-of-inertia parameter
$K:=\frac{4I}{ma^{2}}.$ (128)
Let the precollision states be $(\bm{v},\bm{\omega})$ and $(\bm{v}_{1},\bm{\omega}_{1})$, and let $\bm{k}$ be the unit vector along the line of centers at contact. Define the relative translational velocity
$\bm{g}:=\bm{v}_{1}-\bm{v},$ (129)
and the spin sum and difference
$\bm{\Omega}_{+}:=\bm{\omega}+\bm{\omega}_{1},\qquad\bm{\Omega}_{-}:=\bm{\omega}-\bm{\omega}_{1}.$ (130)
A convenient form of the collision map is
$\bm{v}^{\prime}$ $=\bm{v}+\bm{M},$ $\bm{v}_{1}^{\prime}$ $=\bm{v}_{1}-\bm{M},$ (131)
$\bm{\omega}^{\prime}$ $=\bm{\omega}-\bm{N},$ $\bm{\omega}_{1}^{\prime}$ $=\bm{\omega}_{1}-\bm{N},$ (132)
with
$\bm{M}$ $=\frac{K}{K+1}\left[\bm{g}-\frac{a}{2}\bm{k}\times\bm{\Omega}_{+}+\frac{1}{K}(\bm{k}\cdot\bm{g})\bm{k}\right],$ (133)
$\bm{N}$ $=\frac{2}{aK}\bm{k}\times\bm{M}.$ (134)
These formulas encode conservation of linear momentum, total angular momentum, and kinetic energy together with the no-slip condition at contact; see, for example, Refs. [6, 4, 16].
The relative velocity of the contact points has the same normal component as the center-to-center relative velocity,
$\bm{k}\cdot\bm{g}_{c}=\bm{k}\cdot\bm{g},$ (135)
so the collision rate factor is $\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})$. This identity is used repeatedly below.
VII.2 Dilute-gas estimate of the rotational viscosity $\eta_{r}$
VII.2.1 Homogeneous reference state
Consider a spatially homogeneous state with zero mean flow, uniform density, uniform temperature, and a small uniform mean spin $\bm{\omega}_{\mathrm{h}}(t)$. A natural leading-order reference distribution is the shifted Maxwellian
$f_{\mathrm{h}}(\bm{v},\bm{\omega};t)=n\left(\frac{m}{2\pi k_{B}T}\right)^{3/2}\left(\frac{I}{2\pi k_{B}T}\right)^{3/2}\exp\!\left[-\frac{mv^{2}}{2k_{B}T}-\frac{I|\bm{\omega}-\bm{\omega}_{\mathrm{h}}|^{2}}{2k_{B}T}\right].$ (136)
It satisfies
$\int\bm{\omega}f_{\mathrm{h}}\,\mathrm{d}^{3}v\,\mathrm{d}^{3}\omega=n\bm{\omega}_{\mathrm{h}},\qquad\rho J\bm{\omega}_{\mathrm{h}}=nI\bm{\omega}_{\mathrm{h}}.$ (137)
Our goal is to compute the linear relaxation rate of $\bm{\omega}_{\mathrm{h}}$.
VII.2.2 Collisional production of the spin density
In a binary collision the change in the pair spin sum is
$\Delta\bm{\Omega}_{+}:=(\bm{\omega}^{\prime}+\bm{\omega}_{1}^{\prime})-(\bm{\omega}+\bm{\omega}_{1})=-2\bm{N}.$ (138)
Using Eq. (134),
$\Delta\bm{\Omega}_{+}=-\frac{4}{aK}\bm{k}\times\bm{M}.$ (139)
Substituting Eq. (133) and using the vector identity $\bm{k}\times(\bm{k}\times\bm{a})=\bm{k}(\bm{k}\cdot\bm{a})-\bm{a}$ gives
$\Delta\bm{\Omega}_{+}$ $=-\frac{4}{a(K+1)}\bm{k}\times\bm{g}+\frac{2}{K+1}\Bigl[\bm{k}(\bm{k}\cdot\bm{\Omega}_{+})-\bm{\Omega}_{+}\Bigr].$ (140)
The first term does not contribute after the incoming-hemisphere average over $\bm{k}$: for fixed $\bm{g}$, the integral $\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\bm{k}\times\bm{g})$ vanishes by axial symmetry around $\bm{g}$. The second term produces the linear relaxation.
The dilute collisional production of the spin density may be estimated directly from the pair-collision change in the spin sum and symmetrization over the two collision partners. This gives
$\partial_{t}(nI\omega_{\mathrm{h},i})=\frac{a^{2}I}{2}\int\mathrm{d}^{3}v\,\mathrm{d}^{3}v_{1}\,\mathrm{d}^{3}\omega\,\mathrm{d}^{3}\omega_{1}\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})\,f_{\mathrm{h}}f_{\mathrm{h},1}\,\Delta\Omega_{+,i}.$ (141)
Here $f_{\mathrm{h},1}=f_{\mathrm{h}}(\bm{v}_{1},\bm{\omega}_{1};t)$.
VII.2.3 Angular average
For fixed $\bm{g}$, symmetry implies that the hemisphere tensor integral must be of the form
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})k_{i}k_{j}=A(|\bm{g}|)\delta_{ij}+B(|\bm{g}|)\hat{g}_{i}\hat{g}_{j},$ (142)
where $\hat{\bm{g}}=\bm{g}/|\bm{g}|$. The coefficients are found by taking the trace and the projection along $\hat{\bm{g}}$.
The trace gives
$3A+B$ $=\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})$
$=|\bm{g}|\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\pi/2}\cos\theta\sin\theta\,\mathrm{d}\theta=\pi|\bm{g}|.$ (143)
Projecting with $\hat{g}_{i}\hat{g}_{j}$ gives
$A+B$ $=\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\hat{\bm{g}}\cdot\bm{k})^{2}$
$=|\bm{g}|\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\pi/2}\cos^{3}\theta\sin\theta\,\mathrm{d}\theta=\frac{\pi}{2}|\bm{g}|.$ (144)
Solving these two equations gives
$A=B=\frac{\pi}{4}|\bm{g}|,$ (145)
so
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})k_{i}k_{j}=\frac{\pi|\bm{g}|}{4}(\delta_{ij}+\hat{g}_{i}\hat{g}_{j}).$ (146)
Subtracting the identity tensor gives
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{j}-\delta_{ij})=\pi|\bm{g}|\left(-\frac{3}{4}\delta_{ij}+\frac{1}{4}\hat{g}_{i}\hat{g}_{j}\right).$ (147)
For fixed $|\bm{g}|$, an isotropic average over the direction of $\bm{g}$ uses $\left\langle\hat{g}_{i}\hat{g}_{j}\right\rangle_{\hat{\bm{g}}}=\delta_{ij}/3$, giving
$\left\langle\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{j}-\delta_{ij})\right\rangle_{\hat{\bm{g}}}=-\frac{2\pi}{3}|\bm{g}|\,\delta_{ij}.$ (148)
A subsequent Maxwellian average over $|\bm{g}|$ then gives
$\left\langle\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{j}-\delta_{ij})\right\rangle_{0}=-\frac{2\pi}{3}\,\left\langle|\bm{g}|\right\rangle_{0}\,\delta_{ij}.$ (149)
The mean relative speed of two Maxwellian particles at temperature $T$ is
$\left\langle|\bm{g}|\right\rangle_{0}=4\sqrt{\frac{k_{B}T}{\pi m}},$ (150)
with the derivation recorded in Appendix B.
VII.2.4 Relaxation rate and constitutive matching
Using Eq. (140) in Eq. (141), discarding the $\bm{k}\times\bm{g}$ term whose incoming-hemisphere average vanishes by axial symmetry around $\bm{g}$, and using the fact that $\bm{g}$ and $\bm{\Omega}_{+}$ are statistically independent in the reference state, we obtain
$\partial_{t}(nI\omega_{\mathrm{h},i})$ $=\frac{a^{2}I}{2}\cdot\frac{2}{K+1}\int f_{\mathrm{h}}f_{\mathrm{h},1}\left[\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{j}-\delta_{ij})\right]\Omega_{+,j}\mathrm{d}\Gamma\,\mathrm{d}\Gamma_{1}$
$=-\frac{2\pi a^{2}I}{3(K+1)}\left\langle|\bm{g}|\right\rangle_{0}\int f_{\mathrm{h}}f_{\mathrm{h},1}\,\Omega_{+,i}\,\mathrm{d}\Gamma\,\mathrm{d}\Gamma_{1}.$ (151)
Since $\left\langle\bm{\Omega}_{+}\right\rangle=2\bm{\omega}_{\mathrm{h}}$ and $\int f_{\mathrm{h}}f_{\mathrm{h},1}\,\mathrm{d}\Gamma\,\mathrm{d}\Gamma_{1}=n^{2}$, we obtain
$\partial_{t}\bm{\omega}_{\mathrm{h}}=-\nu_{\mathrm{spin}}\bm{\omega}_{\mathrm{h}},\qquad\nu_{\mathrm{spin}}=\frac{16\sqrt{\pi}}{3(K+1)}na^{2}\sqrt{\frac{k_{B}T}{m}}.$ (152)
Now compare this microscopic decay law with the homogeneous zero-vorticity limit of the macroscopic spin equation,
$\rho J\partial_{t}\bm{\omega}_{\mathrm{h}}=-4\eta_{r}\bm{\omega}_{\mathrm{h}}.$ (153)
Matching Eqs. (152) and (153) gives
$\eta_{r}$ $=\frac{\rho J}{4}\nu_{\mathrm{spin}}=\frac{nI}{4}\nu_{\mathrm{spin}}$
$=\frac{4\sqrt{\pi}}{3(K+1)}n^{2}Ia^{2}\sqrt{\frac{k_{B}T}{m}}=\frac{\sqrt{\pi}K}{3(K+1)}n^{2}ma^{4}\sqrt{\frac{k_{B}T}{m}}.$ (154)
This is positive for every $K>0$, vanishes with $K\to 0$, and scales as $n^{2}$, which is the characteristic signature of a collisional exchange coefficient.
VII.3 Dilute-gas estimate of the transverse spin-diffusion combination $\beta+\gamma$
VII.3.1 Setup and first-order ansatz
We next estimate the transverse combination $\beta+\gamma$. Consider a state with zero mean flow, uniform density, uniform temperature, and a slowly varying transverse mean spin,
$\bm{u}=\bm{0},\qquad n=\mathrm{const.},\qquad T=\mathrm{const.},\qquad\bm{\omega}_{0}(y)=\omega_{0z}(y)\,\bm{e}_{z}.$ (155)
Then $\bm{\nabla}\!\cdot\bm{\omega}_{0}=0$ and the constitutive law reduces to
$m_{zy}=(\beta+\gamma)\,\partial_{y}\omega_{0z}.$ (156)
A local Maxwellian reference state is
$f^{(0)}(y,\bm{c},\bm{\omega})=n\left(\frac{m}{2\pi k_{B}T}\right)^{3/2}\left(\frac{I}{2\pi k_{B}T}\right)^{3/2}\exp\!\left[-\frac{mc^{2}}{2k_{B}T}-\frac{I|\bm{\omega}-\bm{\omega}_{0}(y)|^{2}}{2k_{B}T}\right].$ (157)
Define
$\Omega_{i}:=\omega_{i}-\omega_{0i}(y),\qquad g_{s}:=\partial_{y}\omega_{0z}.$ (158)
Differentiating Eq. (157) with respect to $y$ gives
$\partial_{y}f^{(0)}=\frac{I}{k_{B}T}\Omega_{z}f^{(0)}g_{s},\qquad c_{y}\partial_{y}f^{(0)}=\frac{I}{k_{B}T}c_{y}\Omega_{z}f^{(0)}g_{s}.$ (159)
Hence the relevant microscopic mode is
$X_{zy}:=\Omega_{z}c_{y}.$ (160)
In a leading transport-relaxation or first-Sonine approximation we write
$f^{(1)}_{\mathrm{sd}}=A_{m}X_{zy}f^{(0)}g_{s}.$ (161)
If $\nu_{m}$ denotes the relaxation rate of the mode $X_{zy}$, then projection of the linearized equation gives
$A_{m}=-\frac{I}{k_{B}T\,\nu_{m}}.$ (162)
The corresponding couple stress is
$m_{zy}$ $=-I\int\Omega_{z}c_{y}f^{(1)}_{\mathrm{sd}}\,\mathrm{d}^{3}c\,\mathrm{d}^{3}\omega$
$=-IA_{m}g_{s}\int\Omega_{z}^{2}c_{y}^{2}f^{(0)}\,\mathrm{d}^{3}c\,\mathrm{d}^{3}\omega$
$=-IA_{m}g_{s}n\left\langle\Omega_{z}^{2}\right\rangle\left\langle c_{y}^{2}\right\rangle$
$=-IA_{m}g_{s}n\left(\frac{k_{B}T}{I}\right)\left(\frac{k_{B}T}{m}\right)$
$=\frac{nIk_{B}T}{m\nu_{m}}g_{s}.$ (163)
Matching with Eq. (156) gives
$\beta+\gamma=\frac{nIk_{B}T}{m\nu_{m}}.$ (164)
The task is therefore reduced to computing $\nu_{m}$.
The relaxation calculation below concerns the single transverse component $X_{zy}$. By isotropy, one may equivalently evaluate any fixed Cartesian component pair of the same type, and we may take $(i,j)=(z,y)$ throughout. For notational compactness we nevertheless keep the symbols $i$ and $j$ in the pair-mode algebra, but from this point onward they are fixed component labels rather than dummy indices: Einstein summation is not being used for $i$ or $j$ (that is, there is no sum over $i$ or $j$). Only explicitly repeated auxiliary indices such as $\ell$ are summed.
VII.3.2 Pair mode and collision kinematics
Define the pair mode for such a fixed component pair
$Y_{ij}:=\Omega_{i}c_{j}+\Omega_{1i}c_{1j}.$ (165)
Introduce the sum/difference variables
$\bm{V}:=\bm{c}+\bm{c}_{1},\qquad\bm{g}:=\bm{c}_{1}-\bm{c},\qquad\bm{\Omega}_{\pm}:=\bm{\Omega}\pm\bm{\Omega}_{1}.$ (166)
Then
$Y_{ij}=\frac{1}{2}\Omega_{+,i}V_{j}-\frac{1}{2}\Omega_{-,i}g_{j}.$ (167)
For the rough collision rule, $\bm{V}$ and $\bm{\Omega}_{-}$ are invariants, while
$\Delta\bm{\Omega}_{+}$ $=-\frac{4}{a(K+1)}\bm{k}\times\bm{g}+\frac{2}{K+1}\Bigl[\bm{k}(\bm{k}\cdot\bm{\Omega}_{+})-\bm{\Omega}_{+}\Bigr],$ (168)
$\Delta\bm{g}$ $=-\frac{2K}{K+1}\bm{g}+\frac{aK}{K+1}\bm{k}\times\bm{\Omega}_{+}-\frac{2}{K+1}(\bm{k}\cdot\bm{g})\bm{k}.$ (169)
Equation (169) follows directly from $\bm{g}^{\prime}=\bm{v}_{1}^{\prime}-\bm{v}^{\prime}=\bm{g}-2\bm{M}$ and Eq. (133). Therefore
$\Delta Y_{ij}=\frac{1}{2}(\Delta\Omega_{+,i})V_{j}-\frac{1}{2}\Omega_{-,i}(\Delta g_{j}).$ (170)
In the homogeneous Maxwellian reference state with zero mean spin, the random vectors $\bm{V}$, $\bm{g}$, $\bm{\Omega}_{+}$, and $\bm{\Omega}_{-}$ are mutually independent and satisfy
$\left\langle V_{j}^{2}\right\rangle=\left\langle g_{j}^{2}\right\rangle=2v_{T}^{2},\qquad\left\langle\Omega_{+,i}^{2}\right\rangle=\left\langle\Omega_{-,i}^{2}\right\rangle=2\omega_{T}^{2},$ (171)
with
$v_{T}^{2}:=\frac{k_{B}T}{m},\qquad\omega_{T}^{2}:=\frac{k_{B}T}{I}.$ (172)
Using Eq. (167),
$\left\langle Y_{ij}^{2}\right\rangle$ $=\frac{1}{4}\left\langle\Omega_{+,i}^{2}\right\rangle\left\langle V_{j}^{2}\right\rangle+\frac{1}{4}\left\langle\Omega_{-,i}^{2}\right\rangle\left\langle g_{j}^{2}\right\rangle$
$=\frac{1}{4}(2\omega_{T}^{2})(2v_{T}^{2})+\frac{1}{4}(2\omega_{T}^{2})(2v_{T}^{2})=2v_{T}^{2}\omega_{T}^{2}.$ (173)
VII.3.3 Relaxation rate as a collision bracket
The linear relaxation rate of this fixed-component mode is defined by
$\nu_{m}=-a^{2}n\,\frac{\displaystyle\left\langle\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})\,\Delta Y_{ij}\,Y_{ij}\right\rangle_{0}}{\left\langle Y_{ij}^{2}\right\rangle_{0}}.$ (174)
To evaluate the numerator for this fixed component pair, introduce
$A_{ij}:=\frac{1}{2}\Omega_{+,i}V_{j},\qquad B_{ij}:=-\frac{1}{2}\Omega_{-,i}g_{j},$ (175)
so that $Y_{ij}=A_{ij}+B_{ij}$. Then
$\Delta Y_{ij}Y_{ij}=(\Delta A_{ij})A_{ij}+(\Delta B_{ij})B_{ij}+(\Delta A_{ij})B_{ij}+(\Delta B_{ij})A_{ij}.$ (176)
The mixed terms vanish after averaging because they involve independent zero-mean variables. Thus the numerator splits into two sectors.
Sector A: the $\Omega_{+}V$ contribution.
Using Eq. (170),
$(\Delta A_{ij})A_{ij}=\frac{1}{4}V_{j}^{2}(\Delta\Omega_{+,i})\Omega_{+,i}.$ (177)
The term proportional to $\bm{k}\times\bm{g}$ in Eq. (168) does not contribute because its hemisphere average vanishes by axial symmetry around $\bm{g}$. Therefore,
$\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\Delta A_{ij})A_{ij}\right\rangle_{0}$ $=\frac{1}{4}\left\langle V_{j}^{2}\right\rangle_{0}\frac{2}{K+1}\sum_{\ell=1}^{3}\left\langle\Omega_{+,i}\Omega_{+,\ell}\right\rangle_{0}\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(k_{i}k_{\ell}-\delta_{i\ell})\right\rangle_{0}$
$=\frac{1}{4}(2v_{T}^{2})\frac{2}{K+1}(2\omega_{T}^{2})\left(-\frac{2\pi}{3}\left\langle|\bm{g}|\right\rangle_{0}\right)$
$=-\frac{16\sqrt{\pi}}{3(K+1)}v_{T}^{3}\omega_{T}^{2}.$ (178)
In the second line we used the Maxwellian-averaged tensor identity (149) together with Eq. (150).
Sector B: the $\Omega_{-}g$ contribution.
Here
$(\Delta B_{ij})B_{ij}=\frac{1}{4}\Omega_{-,i}^{2}(\Delta g_{j})g_{j}.$ (179)
From Eq. (169),
$(\Delta g_{j})g_{j}=-\frac{2K}{K+1}g_{j}^{2}+\frac{aK}{K+1}(\bm{k}\times\bm{\Omega}_{+})_{j}g_{j}-\frac{2}{K+1}(\bm{k}\cdot\bm{g})k_{j}g_{j}.$ (180)
The term involving $\bm{\Omega}_{+}$ averages to zero because $\left\langle\bm{\Omega}_{+}\right\rangle=0$. Thus
$(\Delta g_{j})g_{j}\to-\frac{2K}{K+1}g_{j}^{2}-\frac{2}{K+1}(\bm{k}\cdot\bm{g})k_{j}g_{j}.$ (181)
The needed angular integrals are
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})$ $=\pi|\bm{g}|,$ (182)
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})^{2}k_{j}$ $=\frac{\pi}{2}|\bm{g}|g_{j},$ (183)
with the derivation recorded in Appendix A. Using Eq. (181) we obtain
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\Delta g_{j})g_{j}$ $=-\frac{2K}{K+1}\pi|\bm{g}|g_{j}^{2}-\frac{2}{K+1}\cdot\frac{\pi}{2}|\bm{g}|g_{j}^{2}$
$=-\pi\frac{2K+1}{K+1}|\bm{g}|g_{j}^{2}.$ (184)
Therefore
$\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})(\Delta B_{ij})B_{ij}\right\rangle_{0}$ $=\frac{1}{4}\left\langle\Omega_{-,i}^{2}\right\rangle_{0}\left[-\pi\frac{2K+1}{K+1}\left\langle|\bm{g}|g_{j}^{2}\right\rangle_{0}\right]$
$=\frac{1}{4}(2\omega_{T}^{2})\left[-\pi\frac{2K+1}{K+1}\cdot\frac{32}{3\sqrt{\pi}}v_{T}^{3}\right]$
$=-\frac{16\sqrt{\pi}}{3}\frac{2K+1}{K+1}v_{T}^{3}\omega_{T}^{2}.$ (185)
Here we used the Maxwellian moment
$\left\langle|\bm{g}|g_{j}^{2}\right\rangle_{0}=\frac{1}{3}\left\langle|\bm{g}|^{3}\right\rangle_{0}=\frac{32}{3\sqrt{\pi}}v_{T}^{3},$ (186)
proved in Appendix B.
Total rate and coefficient.
Adding Eqs. (178) and (185) gives
$\left\langle\int\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})\Delta Y_{ij}Y_{ij}\right\rangle_{0}=-\frac{32\sqrt{\pi}}{3}v_{T}^{3}\omega_{T}^{2}.$ (187)
Remarkably, the explicit $K$ dependence cancels. Substituting Eqs. (173) and (187) into Eq. (174) gives
$\nu_{m}=\frac{16\sqrt{\pi}}{3}na^{2}\sqrt{\frac{k_{B}T}{m}}.$ (188)
Finally, Eq. (164) yields
$\beta+\gamma$ $=\frac{nIk_{B}T}{m\nu_{m}}$
$=\frac{3I}{16\sqrt{\pi}a^{2}}\sqrt{\frac{k_{B}T}{m}}=\frac{3K}{20}\eta_{0}a^{2},$ (189)
where
$\eta_{0}:=\frac{5}{16a^{2}}\sqrt{\frac{mk_{B}T}{\pi}}$ (190)
is the dilute smooth-hard-sphere shear viscosity.
VII.4 Brief note on the symmetric stress sector
The symmetric stress sector for perfectly rough spheres is classical and may be imported from the first-Sonine Pidduck calculation [20, 4, 16]. In the perfectly rough elastic limit one may write
$\eta=\eta_{0}\frac{6(1+K)^{2}}{6+13K},\qquad\xi=\eta_{0}\frac{(1+K)^{2}}{10K}.$ (191)
These formulas are not rederived here because the algebra is standard and the main purpose of the present paper is to spell out the less familiar antisymmetric and spin-diffusion channels.
VIII Targeted event-driven molecular-dynamics checks
The rough-sphere estimates obtained above can now be confronted directly with a denser set of numerical simulations of the same perfectly rough elastic hard spheres. The goal here remains intentionally limited: we do not attempt a full transport survey, but we do test more stringently the two signatures emphasized by Eqs. (154) and (189). Specifically, the expanded homogeneous dataset is used to assess the predicted collisional $n^{2}$ scaling and roughness dependence of $\eta_{r}$, while the linear finite-$k$ transverse retained-spin response associated with $\beta+\gamma$ is used as a qualitative diagnostic. All runs in this section use event-driven molecular dynamics in a periodic cubic box with $N=8192$ particles and $32$ statistically independent seeds per batch.
VIII.1 Protocol and observables
For the homogeneous runs we prepare states with zero mean flow and a spatially uniform mean spin along $z$. The ensemble-averaged spin signal $\overline{\omega}_{z}(t)$ is fitted by a single exponential,
$\overline{\omega}_{z}(t)\approx\overline{\omega}_{z}(0)e^{-\nu_{\mathrm{spin}}^{\mathrm{EDMD}}t},\qquad\eta_{r}^{\mathrm{EDMD}}=\frac{nI}{4}\nu_{\mathrm{spin}}^{\mathrm{EDMD}},$ (192)
where the second relation is the macroscopic matching condition already used in Eq. (153). For the finite-$k$ runs we initialize a single transverse retained-spin mode and record the complex Fourier amplitudes $\hat{\omega}_{z}(t)$ and $\hat{\zeta}_{z}(t)$. These are fitted to the linearized two-field system
$\partial_{t}\begin{pmatrix}\hat{\zeta}_{z}\\ \hat{\omega}_{z}\end{pmatrix}=\begin{pmatrix}-\dfrac{(\eta+\eta_{r})k^{2}}{\rho}&\dfrac{2\eta_{r}k^{2}}{\rho}\\[8.00003pt] \dfrac{2\eta_{r}}{\rho J}&-\dfrac{(\beta+\gamma)k^{2}+4\eta_{r}}{\rho J}\end{pmatrix}\begin{pmatrix}\hat{\zeta}_{z}\\ \hat{\omega}_{z}\end{pmatrix}$, (193)
which implies the internal consistency condition $A_{12}=Jk^{2}A_{21}$ for the off-diagonal entries of the fitted generator. In the updated homogeneous dataset, we consider a baseline A1 run at $(\phi,K)=(0.020,0.400)$, a density sweep A2 at fixed $K=0.400$ over $0.005\leq\phi\leq 0.050$, and a roughness sweep A3 at fixed $\phi=0.020$ over $0.050\leq K\leq 1.000$. Throughout this section the error bars shown in the figures are bootstrap $16\%$$84\%$ intervals obtained from the seed ensemble.
VIII.2 Homogeneous spin relaxation and the rotational viscosity
Figure 1 shows the baseline homogeneous-spin run at $\phi=0.020$ and $K=0.400$. Over the shaded fit interval the ensemble mean is well described by a single exponential, giving $\nu_{\mathrm{spin}}^{\mathrm{EDMD}}=0.38165$, $\eta_{r}^{\mathrm{EDMD}}=5.21\times 10^{-4}$, and $R^{2}_{\log}=0.9817$. This is the cleanest numerical check in the paper because it targets precisely the homogeneous relaxation mechanism from which Eq. (154) was inferred.
Figure 1: Homogeneous spin relaxation in EDMD for perfectly rough elastic hard spheres at $N=8192$, $\phi=0.020$, and $K=0.400$. The solid curve is the ensemble-averaged magnitude of the mean spin, and the dashed line is the single-exponential fit used to extract $\nu_{\mathrm{spin}}$ and hence $\eta_{r}$ via Eq. (192). The shaded interval marks the fit window.
The density sweep at fixed $K=0.400$ is shown in Fig. 2. The expanded sweep now spans $0.005\leq\phi\leq 0.050$, i.e. a full two decades in $n^{2}$. Over the low- to intermediate-density part of the set the extracted $\eta_{r}$ follows the collisional $n^{2}$ guide closely, rising from $3.04\times 10^{-5}$ at $\phi=0.005$ to $1.27\times 10^{-3}$ at $\phi=0.030$, while the single-exponential quality remains high ($R^{2}_{\log}\approx 0.95$ or better for most points up to $\phi\approx 0.03$). At higher densities the data still grow overall but become visibly more scattered and less perfectly monotone; the fit quality drops to $R^{2}_{\log}=0.8945$ at $\phi=0.035$, $0.7378$ at $\phi=0.0475$, and $0.5900$ at $\phi=0.050$. We therefore interpret Fig. 2 as strong support for the dilute-to-moderate-density $n^{2}$ trend, together with a clear indication that the highest-density points are already feeling departures from a simple single-exponential relaxation picture.
Figure 2: Rotational viscosity extracted from homogeneous-spin EDMD runs as a function of $n^{2}$ at fixed $K=0.400$ for the expanded density sweep $0.005\leq\phi\leq 0.050$. The dashed line is a guide proportional to $n^{2}$, normalized through the better-conditioned points ($R^{2}_{\log}\geq 0.90$).
[…]
Figure 3: Rotational viscosity extracted from homogeneous-spin EDMD runs as a function of the reduced moment-of-inertia parameter $K$ at fixed $\phi=0.020$ for the expanded roughness sweep $0.050\leq K\leq 1.000$. The dashed curve is a guide proportional to $K/(K+1)$, normalized through the better-conditioned points ($R^{2}_{\log}\geq 0.90$).
Taken together, Figs. 13 now provide a substantially broader homogeneous benchmark for the dilute-gas rotational-viscosity estimate. The density sweep covers a full two-decade range in $n^{2}$, and the roughness sweep resolves the interval from weak to order-one roughness. Within the better-conditioned subset of the data, both central signatures of Eq. (154)—collisional $n^{2}$ scaling and growth with roughness—are clearly visible. The deviations concentrated at the highest densities and the smallest $K$ are better interpreted as limits of the simple fitting ansatz than as contradictions of the underlying dilute-gas picture.
VIII.3 Finite-$k$ transverse retained-spin decay
The interpretation of the finite-$k$ runs is essentially unchanged by the present update. Figure 4 shows a representative finite-$k$ run at $\phi=0.020$, $K=0.400$, and $k=0.14996$. Two independent batches (B1 and B2) produced the same qualitative picture, and Fig. 4 shows the cleaner B2 realization. The $\hat{\zeta}_{z}$ channel is captured reasonably well by the linear two-field fit, with $R^{2}_{\zeta}=0.9017$, whereas the $\hat{\omega}_{z}$ channel remains noisy and is not well described by a single fitted generator ($R^{2}_{\omega}=-0.0026$ in the plotted case). Correspondingly, the two off-diagonal estimates of the rotational viscosity do not agree: $\eta_{r}^{(A_{12})}=-1.60\times 10^{-2}$ and $\eta_{r}^{(A_{21})}=9.79\times 10^{-4}$, giving a relative cross-consistency measure of $2.26$. The companion B1 batch yields the same qualitative outcome, with $R^{2}_{\zeta}=0.9156$, $R^{2}_{\omega}=0.0021$, and similarly incompatible off-diagonal $\eta_{r}$ estimates.
Figure 4: Representative finite-$k$ retained-spin EDMD run at $N=8192$, $\phi=0.020$, $K=0.400$, and $k=0.14996$ (batch B2). The solid curves are the ensemble-averaged complex Fourier amplitudes and the dashed curves are the best-fit linear two-field model from Eq. (193). The $\hat{\zeta}_{z}$ component is reproduced reasonably well, but the $\hat{\omega}_{z}$ component remains noisy and the off-diagonal extractions of $\eta_{r}$ are not mutually consistent. In the present data this figure should therefore be interpreted as a qualitative diagnostic rather than as a high-precision coefficient measurement.
The finite-$k$ EDMD data are therefore best viewed as a diagnostic baseline for a more refined transverse-response study. They support the existence of a coupled retained-spin/vorticity relaxation channel, but they do not yet justify a precise coefficient-level claim for $\beta+\gamma$ or for the off-diagonal micropolar coupling. In the context of the present paper that division of labor is still useful: the enlarged homogeneous benchmark now gives a substantially firmer check on the rough-sphere prediction for $\eta_{r}$, while the finite-$k$ runs identify where a future, more accurate numerical study of $\beta+\gamma$ and the off-diagonal coupling will have to improve.
IX Discussion and conclusion
We summarize the main logical points of the derivation.
First, the one-particle balance identities follow directly from the Boltzmann–Curtiss equation once the collision invariants are specified, while for finite-size rough particles the exact local intrinsic-spin balance must keep the antisymmetric-stress torque explicit. The intrinsic-spin balance is most transparent when derived from total angular momentum and then reduced by subtracting the orbital part.
Second, when the mean spin $\bm{\omega}_{0}$ is retained explicitly, the relevant first-order Chapman–Enskog calculation is a generalized one in the sense of extended hydrodynamics. In the bookkeeping adopted here, the residual axial relaxation of the retained-spin manifold is placed at $O(\varepsilon)$ within the collisional-transfer channel, so the retained-spin reference manifold is quasi-equilibrium rather than a strictly instantaneous elimination of spin.
Third, the irreducible decomposition of the first-order source cleanly separates the shear, bulk, spin-diffusion, and axial mismatch channels. The symmetric one-particle kinetic stress contains no axial part, and the spin-gradient sectors are excluded from it by parity. As a result, the rotational viscosity $\eta_{r}$ belongs to the collisional-transfer stress torque channel.
Fourth, the first-order constitutive structure recovers the standard micropolar form,
$\rho\frac{\mathrm{D}\bm{u}}{\mathrm{D}t}$ $=-\bm{\nabla}P+(\eta+\eta_{r})\bm{\nabla}^{2}\bm{u}+\left(\frac{\eta}{3}+\xi-\eta_{r}\right)\bm{\nabla}(\bm{\nabla}\!\cdot\bm{u})+2\eta_{r}\bm{\nabla}\!\times\bm{\omega}_{0}+\rho\bm{F},$
$\rho J\frac{\mathrm{D}\bm{\omega}_{0}}{\mathrm{D}t}$ $=(\beta+\gamma)\bm{\nabla}^{2}\bm{\omega}_{0}+(\alpha+\beta-\gamma)\bm{\nabla}(\bm{\nabla}\!\cdot\bm{\omega}_{0})+2\eta_{r}(\bm{\zeta}-2\bm{\omega}_{0})+\rho\bm{G}.$
The derivation displayed here makes it easy to see where each operator comes from.
Finally, for perfectly rough elastic hard spheres, the dilute-gas estimates (154) and (189) provide explicit physical scales for the antisymmetric stress and the transverse couple stress. The rotational viscosity scales as $n^{2}$ because it is collisional, whereas $\beta+\gamma$ has the scale of an ordinary kinetic transport coefficient. The expanded EDMD checks of Sec. VIII now provide a broader homogeneous benchmark for $\eta_{r}$: the denser A2 density sweep supports the predicted $n^{2}$ scaling over a two-decade range in $n^{2}$, and the A3 roughness sweep makes the expected $K/(K+1)$ trend visible across most of the well-conditioned dataset. At the same time, the highest-density homogeneous points and the finite-$k$ transverse extraction remain numerically more delicate, so the transverse channel should still be interpreted qualitatively rather than as a precision coefficient measurement. Complementary response-theoretic consequences of the retained-spin closure, including EDMD observability and model-discrimination tests, are treated separately in a companion manuscript [25]. What remains open is the complete coefficient-level evaluation of the full axial collisional transfer bracket and of the longitudinal combination $\alpha+\beta-\gamma$ for a concrete microscopic collision operator. That is a natural next step, but it is logically separate from the structural first-order derivation given here.
Appendix A Hemisphere integrals used in the rough-sphere calculation
In this appendix we derive the hemisphere integrals used in the rough-sphere estimates.
Choose coordinates so that $\hat{\bm{g}}=\bm{e}_{3}$ and write
$\bm{k}=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta),\qquad 0\leq\theta\leq\frac{\pi}{2},\qquad 0\leq\phi<2\pi.$ (194)
Then $\bm{k}\cdot\bm{g}=|\bm{g}|\cos\theta$ on the incoming hemisphere. Therefore
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})$ $=|\bm{g}|\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\pi/2}\cos\theta\sin\theta\,\mathrm{d}\theta$
$=\pi|\bm{g}|.$ (195)
Next,
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})^{2}k_{j}$ $=|\bm{g}|^{2}\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\pi/2}\cos^{2}\theta\,k_{j}\sin\theta\,\mathrm{d}\theta.$ (196)
By azimuthal symmetry only the component along $\hat{\bm{g}}$ survives. For $j=3$,
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})^{2}k_{3}$ $=|\bm{g}|^{2}\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\pi/2}\cos^{3}\theta\sin\theta\,\mathrm{d}\theta$
$=\frac{\pi}{2}|\bm{g}|^{2}.$ (197)
Since $g_{j}=|\bm{g}|\hat{g}_{j}$, this is equivalent to
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})^{2}k_{j}=\frac{\pi}{2}|\bm{g}|g_{j}.$ (198)
This is Eq. (183).
The tensor integral (146) was derived in the main text by symmetry plus two contractions. The same result can be checked directly in the chosen coordinates. For example, the $33$ component is
$\int_{\mathbb{S}^{2}}\mathrm{d}\bm{k}\;\Theta(\bm{k}\cdot\bm{g})(\bm{k}\cdot\bm{g})k_{3}^{2}$ $=|\bm{g}|\int_{0}^{2\pi}\mathrm{d}\phi\int_{0}^{\pi/2}\cos^{3}\theta\sin\theta\,\mathrm{d}\theta$
$=\frac{\pi}{2}|\bm{g}|,$ (199)
while the $11$ and $22$ components are equal and satisfy
$2I_{11}+I_{33}=\pi|\bm{g}|.$ (200)
Thus $I_{11}=I_{22}=\pi|\bm{g}|/4$, which again gives Eq. (146).
Appendix B Maxwellian moments of relative and sum variables
Let $\bm{c}$ and $\bm{c}_{1}$ be independent centered Maxwellian velocities with variance $k_{B}T/m$ in each Cartesian component. Then
$\bm{g}=\bm{c}_{1}-\bm{c},\qquad\bm{V}=\bm{c}_{1}+\bm{c}$ (201)
are independent centered Gaussians with variance $2k_{B}T/m$ in each component. Hence the probability density of $g:=|\bm{g}|$ is
$p(g)=4\pi g^{2}\left(\frac{m}{4\pi k_{B}T}\right)^{3/2}\exp\!\left(-\frac{mg^{2}}{4k_{B}T}\right),\qquad g\geq 0.$ (202)
The mean relative speed is
$\left\langle g\right\rangle$ $=4\pi\left(\frac{m}{4\pi k_{B}T}\right)^{3/2}\int_{0}^{\infty}g^{3}\exp\!\left(-\frac{mg^{2}}{4k_{B}T}\right)\mathrm{d}g$
$=4\sqrt{\frac{k_{B}T}{\pi m}},$ (203)
which is Eq. (150).
Similarly,
$\left\langle g^{3}\right\rangle$ $=4\pi\left(\frac{m}{4\pi k_{B}T}\right)^{3/2}\int_{0}^{\infty}g^{5}\exp\!\left(-\frac{mg^{2}}{4k_{B}T}\right)\mathrm{d}g$
$=\frac{32}{\sqrt{\pi}}\left(\frac{k_{B}T}{m}\right)^{3/2}=\frac{32}{\sqrt{\pi}}v_{T}^{3}.$ (204)
By isotropy,
$\left\langle g\,g_{j}^{2}\right\rangle=\frac{1}{3}\left\langle g^{3}\right\rangle=\frac{32}{3\sqrt{\pi}}v_{T}^{3},$ (205)
which is Eq. (186).
For the spin variables, if $\bm{\Omega}$ and $\bm{\Omega}_{1}$ are independent centered Gaussians with variance $k_{B}T/I$ in each component, then $\bm{\Omega}_{\pm}=\bm{\Omega}\pm\bm{\Omega}_{1}$ are independent centered Gaussians with variance $2k_{B}T/I$ in each component. This gives the variances quoted in Eq. (171).
Acknowledgements.
This study was supported by JSPS KAKENHI (Grant Number 22K14177) and JST PRESTO (Grant Number JPMJPR23O7).